OpenStudy (anonymous):

What is the solution of 5a + 12 < 2 and 5a - 12 < 3? A. a < -2 B. a < -1.8 C. a < 2.8 D. a < 3

OpenStudy (anonymous):

@InYourHead

OpenStudy (anonymous):

Help me hermano!!!

OpenStudy (anonymous):

I don't know how to do this.

OpenStudy (anonymous):

I don't either. Hold on.

OpenStudy (anonymous):

Okay. I think I've got it.

OpenStudy (anonymous):

You sure you don't wanna ask someone else? Because I'll tell you the truth: Math really isn't my field. I'm not at all confident in this section as other people are.

OpenStudy (anonymous):

I'll tell you how I did it. You can ask someone else just to be sure, if you like.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@CliffSedge

OpenStudy (anonymous):

There are two equations here. 5a + 12 < 2 And..... 5a - 12 < 3 ------------- I changed the < signs into = signs. So.... 5a + 12 = 2 And..... 5a - 12 = 3

OpenStudy (anonymous):

Now, we need to get all of the number onto the LEFT side of the equal sign. So......here are the two new equations, after we do that..... 5a + 12 - 2 = 0 And..... 5a - 12 - 3 = 0

OpenStudy (anonymous):

There's an easier way..

OpenStudy (anonymous):

a=-5 a=3

OpenStudy (anonymous):

We can simplify those equations. You know...make them simpler. So we have this..... 5a +10 = 0 And.... 5a -15 = 0

OpenStudy (anonymous):

NO.

OpenStudy (anonymous):

I mean...no. Not quite.

OpenStudy (anonymous):

Let's look at the first new, simplified equation.... 5a + 10 = 0

OpenStudy (anonymous):

In this case, a = -2

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

ohh right im doing things in my head

OpenStudy (anonymous):

Now, let's look at the second simplified equation. 5a - 15 = 0 In THIS case....a = -3

OpenStudy (anonymous):

I mean 3.

OpenStudy (anonymous):

Sorry. In this case....a = 3

OpenStudy (anonymous):

So these are the two numbers we have, for a..... a = -2 a = 3

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Now, let's change the = signs BACK into their original forms. We change them back into < signs.

OpenStudy (anonymous):

a < -2 a < 3

OpenStudy (anonymous):

Now, which one is more specific, in this case? We know that ( a ) has to be < 3.

OpenStudy (anonymous):

But we know that ( a ) ALSO has to be < -2

OpenStudy (anonymous):

Ok.

OpenStudy (anonymous):

So you see.....a < -2 is a more SPECIFIC rule.

OpenStudy (anonymous):

Is that tricky to imagine?

OpenStudy (anonymous):

kinda yes.

OpenStudy (anonymous):

Okay. Well, let's look at the first rule that we discovered. a < 3. If this is the only rule, then ( a ) could be anything less than 3. ( a ) could be 2, or 1, or 0....right?

OpenStudy (anonymous):

ohh yes!

OpenStudy (anonymous):

BUT NOW....look at the OTHER rule.... a < -2

OpenStudy (anonymous):

We have TWO rules, not just one. a < 3....AND....a < -2

OpenStudy (anonymous):

( a ) can only be a number that fits BOTH rules.

OpenStudy (anonymous):

Could ( a ) be 1 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

No, it can't. Because 1 IS less than 3. But 1 is NOT less than -2

OpenStudy (anonymous):

a < 3 AND a < -2

OpenStudy (anonymous):

wait soory a can be anything -3 -4 -5

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

and soo on

OpenStudy (anonymous):

Right. So, the answer that I got for this problem was choice A. a < -2

OpenStudy (anonymous):

Because if (a < -2) ..... then it would fit both rules.

OpenStudy (anonymous):

Anything less than -2 would make it so that ( a ) fits both rules.

OpenStudy (anonymous):

Thanks man.

OpenStudy (anonymous):

Yeah. That's my opinion. Get someone else's opinion if you like. I'd recommend it. Sorry, math isn't my best area.